# Coning off totally geodesic boundary components of a hyperbolic manifold

### v1

@@ 3. Tailoring warped products to bound curvature from above
Our first goal is to construct a warping function $f$ on an interval $I\subset\mathbb{R}$ such that $f$ vanishes at the left-hand endpoint of $I$, $f$ makes the metric on the cone agree with the hyperbolic metric on $M$ away from a neighborhood of the cone point, and the cone $I\times_{f}N$ is $\operatorname{CAT}(K)$ for some $K<0$, where $N$ is a boundary component of $M$. We will use a criterion of Alexander and Bishop [AB04] to bound the curvature of our warped product from above. They work in a more general setting $B\times_{f}F$ where $B$ need not be an interval, so we simplify their definitions and result here.
@@
[AB04, §2.2] Let $\mathcal{F}K$ denote the family of solutions to the differential equation $g^{\prime\prime}+Kg=0$. An $\mathcal{F}K$-convex function $f:I\to\mathbb{R}$ for $I\subseteq\mathbb{R}$ an interval is one whose restriction to every sub-interval satisfies the differential inequality $f^{\prime\prime}\geq-Kf$ in the barrier sense, meaning if $g\in\mathcal{F}K$ coincides with $f$ at the endpoints of a sub-interval, then $f\leq g$.
@@
[AB04, Theorem 1.1] Let $I=[t_{0},t_{1})\subset\mathbb{R}$ and let $F$ be a complete $\operatorname{CAT}(K_{F})$ space for some $K_{F}\in\mathbb{R}$. Fix some $K\leq 0$ and suppose $f:I\to[0,\infty)$ is $\mathcal{F}K$-convex, with $f^{-1}(0)=\{t_{0}\}$. Suppose further that $f^{\prime}({t_{0}}^{+})^{2}\geq K_{F}$.
Then the warped product $I\times_{f}F$ is $CAT(K)$.
@@
For all $b>0$ and $0<\delta<\sinh(b)$, there exists $t_{0}<b$ and a function $f\colon\thinspace[t_{0},\infty)\to\mathbb{R}$ such that
@@
Choose $t_{0}<b$ so that the line of slope $\delta$ through the point $(0,t_{0})$ intersects $l_{0}$ somewhere on the interval $(t_{0},b)$. Let $l_{1}$ be this second line. By [FM10, Lemma 4.13], there is a smooth function $g$ on $[t_{0},b]$ whose graph is tangent to $l_{0}$ at $(b,\cosh(b))$ and to $l_{1}$ at $(t_{0},b)$, and whose second derivative is bounded below everywhere. In particular we have the following, for some $\mu>0$.
  1. (1)
@@
$f(t):=\begin{cases}g(t)&t\in[t_{0},b]\\ cosh(t)&t\geq b.\end{cases}$
@@
Then $N$ is $\operatorname{CAT}(\kappa)$.
@@
Since the sectional curvature of $N$ is bounded above by 00, so too is the sectional curvature of $\widetilde{N}$. Thus $\widetilde{N}$ is $\operatorname{CAT}(0)$ (cf. Theorem II.1A.6 in [BH99]), which implies it is $\operatorname{CAT}(\kappa)$. The ball $B$ embeds isometrically so $d_{B}(x,y)$ is bounded above by the corresponding distance in the $\operatorname{CAT}(\kappa)$ comparison triangle. By (2) this also gives an upper bound for $d_{N}(x,y)$, verifying the $\operatorname{CAT}(\kappa)$ inequality. ∎
@@ 4. Proof of Theorem A
For the rest of the section we fix these constants and also fix
$\delta=\frac{\pi}{c},\quad t_{0}<b,\quad K<0$
and a cone warping function $f\colon\thinspace[t_{0},\infty)\to\mathbb{R}$ with parameters $(b,\delta,K)$. Since $\delta<\sinh(b)$ such a function exists by Lemma 3.4.
@@
Let $N$ be a boundary component of $M$. There exists a function $f$ such that the warped cone [^1] We use the notation $C_{f}(N)$ to distinguish from $C(N)=N\times[0,1]/(N\times\{0\})$, the topological cone.
@@
is $\operatorname{CAT}(K)$ for some $K<0$. Moreover the subset $[b,\operatorname{BW}_{M}(\partial M))\times_{f}N$ is (Riemannian) isometric to $N_{\operatorname{BW}_{M}(\partial M)}(N)\smallsetminus N_{b}(N)$.
@@
$[b,\operatorname{BW}_{M}(\partial M))\times_{f}N\subset C_{f}(N)$
@@
Figure 2. On the left is a collar neighborhood of width less than $\operatorname{BW}_{M}(\partial M)$ of a component $N$ of $\partial M$. On the right is the corresponding cone in $X$. (Note that although the cone angle is small in the cartoon it will be at least $2\pi$.)
@@
so that, for fixed $x$, the map $t\mapsto E_{N}(x,t)$ gives a unit speed geodesic in $M$, orthogonal to $N$ and starting at $x$.
@@
$E_{N}(x,t)\sim(t,x)\in C_{f}(N),\quad N\in\mathcal{N},\ x\in N,\ t\in[b,\operatorname{BW}_{M}(\partial M)).$
@@
The path metric space $X$ defined by (italic-$\heartsuit$4.2) is negatively curved.
@@
Note that because $f(t)=\cosh(t)$ on $[b,\operatorname{BW}_{M}(\partial M))$, the gluings are by (Riemannian) isometries. Moreover each piece gives an open set in $X$ which is either locally $\operatorname{CAT}(-1)$ (in the case of $\overline{M}$) or $\operatorname{CAT}(K)$ for some $K<0$ (in the case of the $C_{f}(N)$ pieces). In particular the space $X$ is negatively curved. ∎
Recall that the space $\widehat{M}$ was defined in Definition 1.1 as a quotient of $M\sqcup(\partial M\times[0,1])$ where $\partial M\times\{1\}$ is identified with $\partial M\subset M$, and each component of $\partial M\times\{0\}$ is collapsed to a point. The space $X$ is defined in Definition 4.2.
@@
There is a homeomorphism
@@
    If $d_{M}(x,\partial M)\geq\operatorname{BW}_{M}(\partial M)$, then $\phi(x)=x$.
@@
    For each boundary component $N$ of $M$, the image of $C_{f}(N)$ in $\widehat{M}$ is the cone $N\times[0,1]/\sim$.
  4. (4)
    There is a diffeomorphism $\xi\colon\thinspace[b,\operatorname{BW}_{M}(\partial M))\to[0,\operatorname{BW}_{M}(\partial M))$ so that, for each component $N$ of $\partial M$, each $x\in N$, and each $t\in[b,\operatorname{BW}_{M}(\partial M)$, we have
    $\phi(E_{N}(x,t))=E_{N}(x,\xi(t)).$
@@
To begin, we choose a diffeomorphism $\xi\colon\thinspace[b,\operatorname{BW}_{M}(\partial M))\to[0,\operatorname{BW}_{M}(\partial M))$. Let $E_{\partial M}(x,t)=E_{N}(x,t)$ if $x$ is in the component $N$ of $\partial M$ and $t\in[0,\operatorname{BW}_{M}(\partial M))$. For the orthogonal projection map $\pi_{t}$ defined in Definition 1.7, and for any such $x,t$, we have
$\pi_{t}(E_{\partial M}(x,t))=x.$
Now we define $h\colon\thinspace\overline{M}\to M$ by using $\xi$ to stretch a regular neighborhood of $\partial\overline{M}$ to fit onto a regular neighborhood of $\partial M$:
$h(x)=\begin{cases}x&x\in M\smallsetminus N_{\operatorname{BW}_{M}(\partial M)}(\partial M)\\ E_{\partial M}(\pi_{t}(x),\xi(t))&x\in E_{\partial M}(\partial M,t),\ t\in[0,\operatorname{BW}_{M}(\partial M))\end{cases}.$
This is a homeomorphism because of assumption (A1). If $\iota_{M}\colon\thinspace M\to\widehat{M}$ is the inclusion, we set
@@
Note that for any boundary component $N$ of $M$, the map $\phi_{M}$ already determines $\phi_{C}(t,x)$ for $t\in[b,\operatorname{BW}_{M}(\partial M))$; we next need to extend $\phi_{C}$ to the rest of the cone. To do this, let $\alpha\colon\thinspace[t_{0},b]\to[0,1]$ be an orientation preserving homeomorphism. For $t\in[t_{0},b]$ and $x\in N$, define
$\phi_{C}(t,x)=(x,\alpha(t)),$
[…]
@@ 5. Proof of Theorem B
In this section we retain the assumptions made at the beginning of the last section, in particular fixing the constants $b,c,\delta,t_{0},K$, and the cone warping function $f$ with parameters $(b,\delta,K)$. We moreover fix is a closed locally convex subset $S$ in $M$ and a $b^{\prime}\in(b,\operatorname{BW}_{M}(\partial M))$ so the additional assumptions of Theorem B all hold.
@@
defined for each boundary component $N$ of $M$. The metric on the union comes from gluing together the path metrics on the pieces.
@@
$Y=(S\smallsetminus N_{b}(\partial M))\cup\left(\bigsqcup_{N\in\mathcal{N}}\{(t,s)\in C_{f}(N):s\in P_{b},t\in[t_{0},b]\}\right).$
Figure 3. A cartoon of $S\subset M$ (on the left) and $Y\subset X$ (on the right). The regions above level $b$ are isometric. Again, the cone angle on the right is pictured as small, but in the construction will be large.
@@
Lemma 4.4 gives a homeomorphism
@@
which sends each point $x$ of $\overline{M}$ in $X$ either to $x$ (if it is outside the $\operatorname{BW}_{M}(\partial M)$–neighborhood of the boundary), or to a point on a shortest path from $x$ to $\partial M$. Points inside the $\operatorname{BW}_{M}(\partial M)$–neighborhood of $\partial M$ which are equidistant from the boundary are sent to points which are equidistant from the boundary.
@@
Our assumption 1 together with Lemma 4.4 will ensure that the homeomorphism $\phi$ sends $Y$ to a set which is isotopic to $\widehat{S}$. Specifically, we use the isotopy $\Phi\colon\thinspace[0,b^{\prime}]\times\partial M\to\partial M$ to define an isotopy $\Psi\colon\thinspace[0,1]\times\widehat{M}\to\widehat{M}$ as follows:
  • On $M\smallsetminus N_{b^{\prime}}(\partial M)$: Define $\Psi_{s}(x)=x$ for all $s\in[0,1]$.
@@
  • On $N_{b^{\prime}}(\partial M)$: Recall that we may identify $N_{b^{\prime}}(\partial M)$ with $[0,b^{\prime})\times_{\cosh t}\partial M$. Using the coordinates from this identification, define
    $\Psi_{s}(t,x)=(t,\Phi_{st}\circ(\Phi_{s\xi^{-1}(t)})^{-1}(x)).$
  • On a cone $N\times[0,1]/(N\times\{0\})$: Define
    $\Psi_{s}(x,t)=(\Phi_{sb}^{-1}(x),t).$
The reader may check that this gives a well-defined isotopy taking $\phi(Y)$ to $\widehat{S}$. ∎
It remains to prove the lemmas. Here is some notation that will be helpful. For a point $x$ in some $C_{f}(N)\subset X$, we can write $x=(t(x),p(x))$ where $t(x)$ is in $[t_{0},\operatorname{BW}_{M}(\partial M))$ and $p(x)$ is in $N$. (For $t(x)\in[b,\operatorname{BW}_{M}(\partial M))$ such a point also corresponds to a point in $\overline{M}$, but this ambiguity will cause no problems.)
@@
Fix $y=(t(y),p(y))$ in $Y\cap C_{f}(N)$ for some boundary component $N$. If $t(y)>b$, then $Y$ is locally convex at $y$ by Lemma 5.1, so we suppose $t(y)\leq b$. Let
$L=\frac{2}{3}(t(y)-t_{0})+t_{0}.$
(This is $\frac{2}{3}$ the distance from the cone point to $y$.) We set

### v2

@@ 3. Tailoring warped products to bound curvature from above
Our first goal is to construct a warping function $f$ on an interval $I\subset\mathbb{R}$ such that $f$ vanishes at the left-hand endpoint of $I$, $f$ makes the metric on the cone agree with the hyperbolic metric on $M$ away from a neighborhood of the cone point, and the cone $I\times_{f}N$ is $\operatorname{CAT}(K)$ for some $K<0$, where $N$ is a boundary component of $M$. We will use a criterion of Alexander and Bishop [AB] to bound the curvature of our warped product from above. They work in a more general setting $B\times_{f}F$ where $B$ need not be an interval, so we simplify their definitions and result here.
@@
[AB, §2.2] Let $\mathcal{F}K$ denote the family of solutions to the differential equation $g^{\prime\prime}+Kg=0$. An $\mathcal{F}K$-convex function $f:I\to\mathbb{R}$ for $I\subseteq\mathbb{R}$ an interval is one whose restriction to every sub-interval satisfies the differential inequality $f^{\prime\prime}\geq-Kf$ in the barrier sense, meaning if $g\in\mathcal{F}K$ coincides with $f$ at the endpoints of a sub-interval, then $f\leq g$.
@@
[AB, Theorem 1.1] Let $I=[t_{0},t_{1})\subset\mathbb{R}$ and let $F$ be a complete $\operatorname{CAT}(K_{F})$ space for some $K_{F}\in\mathbb{R}$. Fix some $K\leq 0$ and suppose $f:I\to[0,\infty)$ is $\mathcal{F}K$-convex, with $f^{-1}(0)=\{t_{0}\}$. Suppose further that $f^{\prime}({t_{0}}^{+})^{2}\geq K_{F}$. [^3] Here $f^{\prime}(t_{0}^{+})$ denotes the right-hand derivative at $t_{0}$.
Then the warped product $I\times_{f}F$ is $\operatorname{CAT}(K)$.
Proof.
Notice first that the hypotheses imply that $f$ is strictly increasing on $I$. For $t\in(t_{0},t_{1})$, let
$L_{t}=[t_{0},t]\times_{f|[t_{0},t]}F.$
Lemma 2.3.(2) implies that $L_{t}$ is convex in $I\times_{f}F$. Moreover, [AB, Theorem 1.1] implies that $L_{t}$ is $\operatorname{CAT}(K)$. Any geodesic triangle lies in $L_{t}$ for some $t<t_{1}$, so the entire warped product is $\operatorname{CAT}(K)$. ∎
@@
Proof.
Let $J\subseteq I$ be a closed interval, and let $g\colon\thinspace J\to\mathbb{R}$ be the function which agrees with $f$ on the endpoints of $J$ and satisfies the differential equation $g^{\prime\prime}+Kg=0$. Let h=fgh=f-g; we must show that $h\leq 0$ on $J$.
Suppose by contradiction that $h(t_{0})>0$ for some $t_{0}\in J$. Shrinking $J$ if necessary, we may suppose that $h$ is positive everywhere on the interior of $J$. Almost everywhere on $J$, we have
$h^{\prime\prime}(t)$ $=f^{\prime\prime}(t)-g^{\prime\prime}(t)$
$\geq-Kf(t)+Kg(t)=-Kh(t)$
$>0.$
Let $\ell$ be the left-hand endpoint of $J$. Since $h(t)>0$ for $t$ in the interior of $J$, we must have $h^{\prime}(\ell)\geq 0$. But then for any $t$ in the interior of $J$, we have $h^{\prime}(t)=h^{\prime}(\ell)+\int_{\ell}^{t}h^{\prime\prime}(y)\ dy>0$. In particular $h$ is strictly increasing on $J$. This contradicts the fact that $h$ vanishes at the right-hand endpoint of $J$. ∎
@@
For all $b>0$ and $0<\delta<\sinh(b)$, there exists $t_{0}<b$, $K<0$, and a function $f\colon\thinspace[t_{0},\infty)\to\mathbb{R}$ such that
@@
Choose $t_{0}<b$ so that the line of slope $\delta$ through the point $(t_{0},0)$ intersects $l_{0}$ somewhere on the interval $(t_{0},b)$. Let $l_{1}$ be this second line. By [FM, Lemma 4.13], there is a smooth function $g$ on $t_{0},b]$ whose graph is tangent to $l_{0}$ at $(b,\cosh(b))$ and to $l_{1}$ at $(t_{0},b)$, and whose second derivative is bounded below everywhere. In particular $g$ has the following properties, for some $\mu>0$.
@@
$f(t):=\begin{cases}g(t)&t\in[t_{0},b]\\ \cosh(t)&t\geq b.\end{cases}$
@@
Then $N$ is (globally) $\operatorname{CAT}(\kappa)$.
@@
Since the sectional curvature of $N$ is bounded above by 0, so too is the sectional curvature of $\widetilde{N}$. Thus $\widetilde{N}$ is $\operatorname{CAT}(0)$ (cf. Theorem II.1A.6 in [BH]), which implies it is $\operatorname{CAT}(\kappa)$ (see [BH, Theorem II.1.12]). The ball $B$ lifts isometrically to $\widetilde{N}$, so $d_{B}(x,y)$ is bounded above by the corresponding distance in the $\operatorname{CAT}(\kappa)$ comparison triangle. By (2) this also gives an upper bound for $d_{N}(x,y)$, verifying the $\operatorname{CAT}(\kappa)$ inequality. ∎
@@ 4. Proof of Theorem A
For the rest of the section we fix these constants, and let $\delta=\frac{\pi}{c}$. Lemma 3.4 tells us there are constants $K<0$, $t_{0}<b$ and and a cone warping function $f\colon\thinspace[t_{0},\infty)\to\mathbb{R}$ with parameters $(b,\delta,K)$. We also fix these numbers $K,t_{0}$ and function $f$.
@@
Let $N$ be a boundary component of $M$. The warped cone
@@
is $\operatorname{CAT}(K)$. [^4] We use the notation $C_{f}(N)$ to distinguish from $C(N)=N\times[0,1]/(N\times\{0\})$, the topological cone. Moreover the subset $b,\operatorname{BW}_{M}(\partial M))\times_{f}N$ is (Riemannian) isometric to $N_{\operatorname{BW}_{M}(\partial M)}(N)\smallsetminus N_{b}(N)$.
@@
$b,\operatorname{BW}_{M}(\partial M))\times_{f}N\subset C_{f}(N)$
@@
Figure 2. In the middle is a collar neighborhood of width less than $\operatorname{BW}_{M}(\partial M)$ of a component $N$ of $\partial M$. $\widehat{M}$ is on the left. On the right is the corresponding cone in $X$, with the removed portion of $M$ in dashed grey.
@@
so that, for fixed $p$, the map $t\mapsto E_{N}(p,t)$ gives a unit speed geodesic in $M$, orthogonal to $N$ and starting at $p$.
@@
$E_{N}(p,t)\sim(t,p)\in C_{f}(N),\quad N\in\mathcal{N},\ p\in N,\ t\in[b,\operatorname{BW}_{M}(\partial M)).$
@@
The path metric space $X$ defined by ($\heartsuit$4.2) is negatively curved.
@@
Recall from Lemma 4.1 that for points $(t,p)$ such that $t\in[b,\operatorname{BW}_{M}(\partial M)$, the gluings are by Riemannian isometries. Moreover each piece gives an open set in $X$ which is either locally $\operatorname{CAT}(-1)$ (in the case of $\overline{M}$) or $\operatorname{CAT}(K)$ for some $K<0$ (in the case of the $C_{f}(N)$ pieces). In particular the space $X$ is negatively curved. ∎
Recall that the space $\widehat{M}$ was defined in Definition 1.1 as a quotient of $M\sqcup(\partial M\times[0,1])$ where $\partial M\times\{1\}$ is identified with $\partial M\subset M$, and each component of $\partial M\times\{0\}$ is collapsed to a point. The space $X$ is defined in Definition 4.2. It will be convenient to refer to points in some $C_{f}(N)$ by their coordinates:
Notation 4.4.
For $N$ a boundary component of $X$, and $\ast$ the cone point of $C_{f}(N)$, we define functions
$t\colon\thinspace C_{f}(N)\to[t_{0},\operatorname{BW}_{M}(\partial M))\quad\mbox{and}\quad p\colon\thinspace C_{f}(N)\smallsetminus\{\ast\}\to N$
by projecting onto coordinates. For $x\in C_{f}(N)$, we write $x=(t(x),p(x))$. (The second coordinate is irrelevant when $t(x)=t_{0}$.)
@@
Let
$\xi\colon\thinspace[b,\operatorname{BW}_{M}(\partial M))\to[0,\operatorname{BW}_{M}(\partial M))$
and
$\alpha\colon\thinspace[0,b]\to[0,1]$
be orientation-preserving homeomorphisms. There is then a homeomorphism
@@
    If $x$ is not in $C_{f}(N)$ for any boundary component $N$ of $M$, then $\phi(x)=x$.
@@
    Let $N$ be a boundary component of $M$, and $y\in C_{f}(N)$. Then, identifying $N_{\operatorname{BW}_{M}(\partial M)}(N)\subset M$ with $0,\operatorname{BW}_{M}(\partial M))\times_{\cosh(t)}N$,
    $\phi(y)=\begin{cases}(p(y),\alpha(t(y)))\in N\times[0,1]/\sim&t(y)\leq b\\ (\xi(t(y)),p(y))\in N_{\operatorname{BW}_{M}(\partial M)}(N)&t(y)\geq b.\end{cases}$
@@
If $p$ is in the component $N$ of $\partial M$ and $t\in[0,\operatorname{BW}_{M}(\partial M))$, we set $E_{\partial M}(p,t)=E_{N}(p,t)$. For the orthogonal projection map $\pi_{t}$ defined in Definition 1.7, and for any such $p,t$, we have
$\pi_{t}(E_{\partial M}(p,t))=p.$
Now we define $h\colon\thinspace\overline{M}\to M$ by using $\xi$ to stretch a regular neighborhood of $\partial\overline{M}$ in $\overline{M}$ to fit onto a regular neighborhood of $\partial M$ in $M$:
$h(x)=\begin{cases}x&x\in M\smallsetminus N_{\operatorname{BW}_{M}(\partial M)}(\partial M)\\ E_{\partial M}(\pi_{t}(x),\xi(t))&x\in E_{\partial M}(\partial M,t),\ t\in[b,\operatorname{BW}_{M}(\partial M))\end{cases}.$
Figure 3. See Figure 2 for descriptions of the objects on the left and right. The pink brackets indicate $\phi_{M}$; the blue brackets indicate $\phi_{C}$.
This is clearly a homeomorphism. If $\iota_{M}\colon\thinspace M\to\widehat{M}$ is the inclusion, we define $\phi_{M}\colon\thinspace\overline{M}\to\widehat{M}$ by
@@
See Figure 3. For any boundary component $N$ of $M$, the map $\phi_{M}$ determines $\phi_{C}(t,p)$ for $t\in[b,\operatorname{BW}_{M}(\partial M))$; it remains to consistently define $\phi_{C}$ on the rest of the cones $C_{f}(N)$.
For $t\in[t_{0},b]$ and $p\in N$, define
(3) $\phi_{C}(t,p)=(p,\alpha(t)).$
Warped product coordinates appear on the left-hand side of (3); the coordinates on the right-hand side come from the description of the topological cone-off in equation ($\ddagger$}1.1) of Definition 1.1. In the definition ($\heartsuit$4.2), the points $E_{N}(p,b)\in\overline{M}$ and $(b,p)\in C_{f}(N)$ are identified in $X$; we need to check these go to the same point in $\widehat{M}$ under $\phi_{C}$ and $\phi_{M}$. Indeed
$\phi_{M}(E_{N}(p,b))=\iota_{M}(h(E_{N}(p,b)))=\iota_{M}(p,\xi(b))=\iota_{M}(p,0)=(p,1).$
We also have $\phi_{C}(b,p)=(p,\alpha(b))=(p,1)$ as desired, so the map $\phi$ assembled from $\phi_{M}$ and the $\phi_{C}$ is well-defined. Since $\phi$ is a continuous bijection of compact metric spaces, it is a homeomorphism. The conditions (1)–(3) are satisfied by construction. ∎
@@ 5. Proof of Theorem B
In this section we retain the assumptions made at the beginning of the last section, in particular fixing the constants $b,c,\delta,t_{0},K$, and the cone warping function $f$ with parameters $(b,\delta,K)$. We moreover fix a closed locally convex subset $S$ in $M$ and a $b^{\prime}\in(b,\operatorname{BW}_{M}(\partial M))$ so the additional assumptions of Theorem B all hold.
@@
defined for each boundary component $N$ of $M$. The metric on the union comes from gluing together the path metrics on the pieces. Recall the notation for coordinate functions $(t(x),p(x))$ for points $x$ in $C_{f}(N)$ outlined in 4.4
@@
$Y=(S\smallsetminus N_{b}(\partial M))\cup\left(\bigsqcup_{N\in\mathcal{N}}\left\{x\in C_{f}(N)\ \mid\ p(x)\in P_{b},\ t(x)\leq b\right\}\right)/\sim.$
Figure 4. A cartoon of $S\subset M$ (on the left) and $Y\subset X$ (on the right). The regions above level $b$ are Riemannian isometric.
@@
We choose a homeomorphism $\xi\colon\thinspace[b,\operatorname{BW}_{M}(\partial M))\to[0,\operatorname{BW}_{M}(\partial(M)))$ so that $\xi$ restricts to the identity on the interval $b^{\prime},\operatorname{BW}_{M}(\partial{M}))$. Lemma 4.5 gives a homeomorphism
@@
which sends each point $x$ of $\overline{M}$ in $X$ either to $x$ (if it is outside the $b^{\prime}$–neighborhood of the boundary), or to a point on a shortest path from $x$ to $\partial M$. Points inside the $\operatorname{BW}_{M}(\partial M)$–neighborhood of $\partial M$ which are equidistant from the boundary are sent to points which are equidistant from the boundary. More precisely, for $y\in C_{f}(N)$ there are two cases. When $t(y)\geq b$, the point $\phi(y)$ lies in $N_{\operatorname{BW}_{M}(\partial{M})}(N)\cong[0,\operatorname{BW}_{M}(\partial{M}))\times_{\cosh(t)}N$. With respect to these warped product coordinates, and using Notation 4.4, we have
$\phi(y)=(\xi(t(y)),p(y)).$
When $t(y)<b$, the point $\phi(y)$ lies in the cone $C(N)=N\times[0,1]/N\times\{0\}$; its coordinates are given by
$\phi(y)=(p(y),\alpha(t(y))).$
@@
Our assumption 1 together with Lemma 4.5 will ensure that the homeomorphism $\phi$ sends $Y$ to a set which is isotopic to $\widehat{S}$. Specifically, we use the isotopy $\Phi\colon\thinspace[0,b^{\prime}]\times\partial M\to\partial M$ to define an isotopy $\Psi\colon\thinspace[0,1]\times\widehat{M}\to\widehat{M}$ piecewise as follows:
  • On $Q_{1}:=M\smallsetminus N_{b^{\prime}}(\partial M)$, define $\Psi_{s}(x)=x$ for all $s\in[0,1]$.
@@
  • On $Q_{2}:=N_{b^{\prime}}(\partial M)$. Recall that we may identify $N_{b^{\prime}}(\partial M)$ with $0,b^{\prime})\times_{\cosh t}\partial M$. Using the coordinates from this identification, define
    (4) $\Psi_{s}(t,p)=(t,\Phi_{st}\circ(\Phi_{s\xi^{-1}(t)})^{-1}(p)).$
  • On the disjoint union $Q_{3}$ of topological cones $N\times[0,1]/(N\times\{0\})$: Define
    $\Psi_{s}(p,\zeta)=(\Phi_{sb}^{-1}(p),\zeta).$
We first verify that $\Psi_{s}$ is a homeomorphism for each $s\in[0,1]$. To check that $\Psi_{s}$ is well-defined, we must see what happens when the $t$ variable in $N_{b^{\prime}}(\partial M)$ is equal to $b^{\prime}$ or to 0. Since $\xi(b^{\prime})=b^{\prime}$, Equation (4) gives
$\Psi_{s}(b^{\prime},p)=(b^{\prime},p),$
and this agrees with the definition on $M\smallsetminus N_{b^{\prime}}(\partial M)$. For $t=0$, Equation (4) gives
$\Psi_{s}(0,p)=(0,\Psi_{sb}^{-1}(p)),$
which agrees with the definition on the topological cone. Now for each piece $Q\in\{M\smallsetminus N_{b^{\prime}}(\partial M),N_{b^{\prime}}(\partial M),\partial N\times[0,1]/\sim\}$, $\Psi_{s}$ maps $Q$ homeomorphically to itself, so $\Psi_{s}$ is a continuous bijection. Since $\widehat{M}$ is compact Hausdorff, this implies that $\Psi_{s}$ is a homeomorphism.
Define $\widetilde{Q}_{i}:=\phi^{-1}(Q_{i})$.
Claim 5.4.
$\Psi_{1}\circ\phi$ takes $\widetilde{Q}_{1}\cap Y$ bijectively to $Q_{1}\cap\widehat{S}$.
Note that $Q_{1}$ and $\widetilde{Q}_{1}$ are the same subset of $M$, and $\Psi_{s}\circ\phi$ is the identity on this subset for any $s$. (This is because $\xi(t)=t$ for all $t\in[b^{\prime},\operatorname{BW}_{M}(\partial M))$.)
Claim 5.5.
$\Psi_{1}\circ\phi$ takes $\widetilde{Q}_{2}\cap Y$ bijectively to $Q_{2}\cap\widehat{S}$
We have $\widetilde{Q}_{2}\cap Y=\{y\in Y:t(y)\in[b,b^{\prime}]\}$ and $Q_{2}\cap\widehat{S}=\{x\in S:d(x,\partial M)\leq b^{\prime}\}$. As we have already shown that $\Psi_{1}$ is a homeomorphism, it suffices to show that the image of the first set is contained in the second, and the preimage of the second is contained in the first. For $y\in\widetilde{Q}_{2}\cap Y$, we have
$\Psi_{1}(\phi(y))=(\xi(t),\Phi_{\xi(t)}\circ\Phi_{t}^{-1}(p(y))).$
Since $p(y)=\pi_{t}(y)\in P_{t}$, we have $p(\Psi_{1}(\phi(y)))=\Phi_{\xi(t)}\circ\Phi_{t}^{-1}(p(y))\in P_{\xi(t)}$. This implies $\Psi_{1}(\phi(y))\in S$. Moreover $\xi$ maps $b,b^{\prime}]$ to $0,b^{\prime}]$ so $d(\Psi_{1}(\phi(y)),\partial M)=\xi(t)\leq b^{\prime}$.
Now let $x\in S\cap Q_{2}$ and write $(t,p)=(t(x),p(x))$. We have
$(\Psi_{1}\circ\phi)^{-1}(x)=(\xi^{-1}(t),\Phi_{\xi^{-1}(t)}\circ\Phi_{t}^{-1}(p)).$
Since $t\leq b^{\prime}$, it follows that $\xi^{-1}(t)\in[b,b^{\prime}]$. Because $x\in S$, we have that $p$ is in $P_{t}$, and so $p((\Psi_{1}\circ\phi)^{-1}(x))=\Phi_{\xi^{-1}(t)}\circ\Phi_{t}^{-1}(p)\in P_{\xi^{-1}(t)}$. Therefore $(\Psi_{1}\circ\phi)^{-1}(x)$ is in $\widetilde{Q}_{2}\cap Y$.
Claim 5.6.
$\Psi_{1}\circ\phi$ takes $\widetilde{Q}_{3}\cap Y=\{y\in Y:t(y)\leq b\}$ bijectively to $Q_{3}\cap\widehat{S}$.
Let $y\in\widetilde{Q}_{3}\cap Y$. For $(p,t)=(p(y),t(y))$ we have
$\Psi_{1}(\phi(y))=(\Phi_{b}^{-1}(p),\alpha(t)).$
In this case $p\in P_{b}$, so $p(\Psi_{1}(\phi(y)))\in P_{0}$ by 2. Additionally, because $t\leq b$, $\alpha(t)\in[0,1]$ is defined. In particular $\Psi_{1}(\phi(y))\in Q_{3}\cap\widehat{S}$.
Now let $(p,\zeta)\in Q_{3}\cap\widehat{S}$. We have
$(\Psi_{1}\circ\phi)^{-1}(p,\zeta)=(\Phi_{b}(p),\alpha^{-1}(\zeta)).$
Note that $\alpha^{-1}(\zeta)\leq b$, and 1 gives $\Phi_{b}(p)\in P_{b}$. Therefore $(\Psi_{1}\circ\phi)^{-1}(p,\zeta)\in\widetilde{Q}_{3}\cap Y$.
Pushing forward the metric from $X$ under $\Psi_{1}\circ\phi$, we obtain a metric with respect to which $\widehat{S}$ is locally convex. (The isotopy $\Psi$ justifies Remark 1.8.) ∎
It remains to prove the lemmas.
@@
Fix $y=(t(y),p(y))$ a non-cone-point in $Y\cap C_{f}(N)$ for some boundary component $N$. If $t(y)>b$, then $Y$ is locally convex at $y$ by Lemma 5.1, so we suppose $t(y)\leq b$. Let
$L=\frac{1}{3}(2t(y)+t_{0}).$
(The set of points so that $t(x)=L$ is the copy of $N$ which is $\frac{2}{3}$ the distance from the cone point to $y$.) We set